# Enhancing Security in Multi-Robot Systems through Co-Observation Planning, Reachability Analysis, and Network Flow

### v1

@@ Abstract
This paper addresses security challenges in multi-robot systems (MRS) where adversaries may compromise robot control, risking unauthorized access to forbidden areas. We propose a novel multi-robot optimal planning algorithm that integrates mutual observations and introduces reachability constraints for enhanced security. This ensures that, even with adversarial movements, compromised robots cannot breach forbidden regions without missing scheduled co-observations. The reachability constraint uses ellipsoidal over-approximation for efficient intersection checking and gradient computation. To enhance system resilience and tackle feasibility challenges, we also introduce sub-teams. These cohesive units replace individual robot assignments along each route, enabling redundant robots to deviate for co-observations across different trajectories, securing multiple sub-teams without requiring modifications. We formulate the cross-trajectory co-observation plan by solving a network flow coverage problem on the checkpoint graph generated from the original unsecured MRS trajectories, providing the same security guarantees against plan-deviation attacks. We demonstrate the effectiveness and robustness of our proposed algorithm, which significantly strengthens the security of multi-robot systems in the face of adversarial threats.
@@ I Introduction
Multi-robot systems (MRS) have found wide applications in various fields. While offering numerous advantages, the distributed nature and dependence on network communication render the MRS vulnerable to cyber threats, such as unauthorized access, malicious attacks, and data manipulation [1]. This paper addresses a specific scenario in which robots are compromised by attacker and directed to forbidden regions, which may contain security-sensitive equipment or human workers. Such countering deliberate deviations, termed plan-deviation attacks, are identified and addressed in previous studies [2, 3, 4, 5, 6]. As a security measure, we utilized onboard sensing capabilities of the robots to perform inter-robot co-observations and check for unusual behavior. These mutual observation establish a co-observation schedule alongside the path, ensuring that any attempts by a compromised robot to violate safety constraints (such as transgressing forbidden regions) would break the observation plan and be promptly detected.
@@
A preliminary version of the paper was presented in [6, 5]. Extanding the grid-world solution from [2] to continuous configuration spaces, we incorporate the co-observation planning as constraints in the ADMM-based trajectory optimization solver to accomodate for more flexible objectives. In this paper, we incorporate additional reachability analysis during the planning phase, incorporating constraints based on sets of locations that agents could potentially reach, referred to as reachability regions, as a novel perspective to the existing literature. This approach enforces an empty intersection between forbidden regions and the reachability region during trajectory optimization, preventing undetected attacks if an adversary gains control of the robots. Inspired by the heuristic sampling domain introduced by [7] within the context of the $\mathtt{RRT^{*}}$ path planning algorithm, we utilize an ellipsoidal boundary to constrain the search space and formulate the ellipsoid as the reachability region constraint. We present a mathematical formulation of reachability regions compatible with the solver in [6] as a spatio-temporal constraint.
@@
However, such plans are not guaranteed to exist, and the secured trajectory always come at the cost of overall system performance. As an extension of the previous works [6, 5], we also address the feasibility and optimality challenges. For a MRS with unsecured optimal trajectorys (without security constraints), we introduce redundant robots and formed into sub-teams with the original ones. These redundant robots are assigned the task of establishing additional co-observations, termed cross-trajectory co-observations, within and across different sub-teams. The proposed algorithm focuses on the movement plans for the additional robots, distinct from those dedicated to task objectives, indicating when they should stay with their current sub-team and when they should deviate to join another. This strategy allows sub-teams to preserve the optimal unsecured trajectories (as illustrated in Figure 1) without requiring the entire sub-team to maintain close proximity to other teams during co-observation events. The cross-trajectory co-observation problem is transformed as a multi-agent path finding problem on roadmap (represented as directed graph) and solved as a network flow problem [8].
@@
Figure 1: 1b Limited by the co-observation requirement, both red and blue robot follow the co-observation secured routes (solid lines) and abandon the optimal ones (dashed line). 1c Through cross-trajectory co-observations, the blue team sends one robot to follow the red team (solid blue line) and performs co-observation while having the rest of the robots following the optimal trajectory.
@@
Related research. Trajectory planning problem of MRS systems remained as a subject of intense study for many decades, and optimization is a common approaches in such area. Optimization based approaches are customizable to a variable of constraints (e.g. speed limit, avoid obstacles) and task specifications (e.g. maximum surveillance coverage, minimal energy cost). In constract to our work in Section II, many motion planning tasks require a non-convex constraint problem formulation, most contributors focused on convex problems, and only allow for a few types of pre-specified non-convex constraints through convexification[9][10][11]. Several optimization techniques like MIQP [12] and ADMM [13] have been used to reduce computational complexity, and to incorporate more complex non-convex constraints.
@@
Graph-based search is another extensively explored approach in trajectory planning problems, such as formation control [14, 15] and Multi-Agent Path Finding (MAPF) problems [16]. In traditional MAPF formulations, environments are abstracted as graphs, with nodes representing positions and edges denoting possible transitions between positions and solved as a network flow problem [8, 17]. This formulation allows for the application of combinatorial network flow algorithms and linear program techniques, offering efficient and more flexible solutions to the planning problem.
In Section II, we incorporate reachability constraints to increase overall security Reachability analysis is a important step in the security and safety verification synthesis for cyber-physical systems (CPSs) [18, 19]. One typical approach involves computing an over-approximation of the reachable space for checking safety properities. Previous studies [20, 21, 22] often use geometric representation, such as zonotopes and ellipsoids, as a compact enclosure of the reachability sets. In the realm of online safety assessment against cyber attacks, [23] computed the reachable set of CPS states achievable by potential cyber attacks and compared it with the safe region, set based on current state estimation and environmental conditions.
Paper contributions. In this paper, two main contributions have been presented.
@@
  • We present an innovative method to integrate reachability analysis into the ADMM-based optimal trajectory solver for multi-robot systems, preventing attackers from executing undetected attacks by simultaneously entering forbidden areas and adhering to co-observation schedules.
@@
  • We introduce additional robots to form sub-teams for both intra-sub-team and cross-sub-team co-observations. A new co-observation planning algorithm is formulated that can generate a resilient multi-robot trajectory with a co-observation plan that still preserve the optimal performance against arbitrary tasks. We also find the minimum redundant robots required for the security.
@@
Paper Outline. The rest of this paper is organized as follows. Section IIprovide an overview of the ADMM-based optimal trajectory solver and introduce the security constraints including co-observation constraints and ellipsoidal reachability constraints. Section III introduced the enhanced security planning algorithm by the incorporation of redundant robots for cross-trajectory co-observation. Section IV concludes this article.
@@ II Secured Multi-robot trajectory planning
We formulate the planning problem as an optimal trajectory optimization problem to minimize arbitrary smooth objective functions. We denote as $q_{ij}\in m$ the position of agent $i$ at the discrete-time index $j$, with $m$ representing the dimension of the state space. For a total of $n_{p}$ robots, and a task time horizon of $T$, a total of $n_{p}$ trajectories can be represented as an aggregated vector $\mathbf{q}\in nmT$. The overall goal is to minimize or maximize an objective function $\varPhi(\mathbf{q})$ under a set of nonlinear constraints described by a set $\Omega$, which is given by the intersection of spatio-temporal sets given by traditional path planning constraintsand the security constraints (co-observation schedule, reachability analysis). Formally:
@@
Robots are tasked to navigate, collecting sensory data to construct a corresponding map for the slowly-varying field, denoted as $\mathbf{x}$, of an unknown environment (see Figure 3b). The map comprises points of interest arranged on a grid, each associated with a slowly-changing value. Our goal is to find paths that minimize uncertainty and effectively reconstruct the field. We associate a Kalman Filter (KF) (cf. [24]) to each point $j$, and use it to track the uncertainty through its estimated covariance $P_{j}$. The updates of the filters are based on measurements from the robot centered at $q$, and we use a Gaussian radial basis function modeling spatially-varying measurement quality. The optimization objective $\varPhi(\mathbf{q})$ is written as the maximum uncertainty $\max_{j}P_{j}$ (detailed formulation can be found in [6]).
To incorporate the security against malicious deviations (through co-observation and reachability analysis) as constraints in the optimal trajectory planning problem, we define the security as:
@@
A multi-robot trajectory plan is secure again plan-deviation attacks if it ensure that any potential deviations to these forbidden regions will cause the corresponding robot to miss their next co-observation with other robots.
@@ II-A1 Differentials
We define the differential of a map $f(x):m\to n$ at a point $x_{0}$ as the unique matrix $\partial_{x}f\in n\times m$ such that
$\left.\frac{\mathrm{d}}{\mathrm{d}t}f\bigl{(}x(t)\bigr{)}\right|_{t=0}=\partial_{x}f\bigl{(}x(0)\bigr{)}\dot{x}(0)$ (2)
@@
where $t\mapsto x(t)\in n$ is a smooth parametric curve such that $x(0)=x$ with any arbitrary tangent $\dot{x}(0)$. For brevity we will use $\dot{f}$ for $\frac{\mathrm{d}}{\mathrm{d}t}f$ and $\partial_{x}f$ for $\frac{\partial f}{\partial x}$.
@@
With a slight abuse of notation, we use the same notation $\partial_{x}f$ for the differential of a matrix-valued function with scalar arguments $f:R\to m\times n$. Note that in this case (2) is still formally correct, although the RHS needs to be interpreted as applying $\partial_{x}f$ as a linear operator to $\dot{x}$.
II-A2 Householder rotations
We define a differentiable transformation to a canonical ellipse that is used in the derivation of the reachability constraints in Sections II-C and II-H. This transformation includes a rotation derived from a modified version of Householder transformations [25]. With respect to the standard definition, our modification ensures that the final operator is a proper rotation (i.e., not a reflection). We call our version of the operator a Householder rotation. In this section we derive Householder rotations and their differentials for the 3-D case; the 2-D case can be easily obtained by embedding it in the $z=0$ plane.
Definition 2
Let $\nu_{\mathcal{F}}$ and $\nu_{\mathcal{E}}$ be two unitary vectors ($\lVert\nu_{\mathcal{F}}\rVert=\lVert\nu_{\mathcal{E}}\rVert=1$). Define the normalized vector $u$ as
$u^{\prime}=\nu_{\mathcal{F}}+\nu_{\mathcal{E}},\quad u=\frac{u^{\prime}}{\lVert u^{\prime}\rVert}.$ (3)
The Householder rotation $H(\nu_{\mathcal{F}},\nu_{\mathcal{E}})$ is defined as
$H(\nu_{\mathcal{F}},\nu_{\mathcal{E}})=2uu^{\mathrm{T}}-I.$ (4)
Here $H$ is a rotation mapping $\hat{\nu_{\mathcal{F}}}$ to $\hat{\nu_{\mathcal{E}}}$, as shown by the following.
Proposition 1
The matrix $H$ has the following properties:
  1. 1.
    It is a rotation, i.e.
    1. (a)
      $H^{\mathrm{T}}H=I$;
    2. (b)
      $\det(H)=1$.
  2. 2.
    $\nu_{\mathcal{E}}=H\nu_{\mathcal{F}}$.
Proof:
See Appendix -A. ∎
We compute the differential of $H$ implicitly using its definition (2). We use the notation $[v]_{\times}\!:3\to 3\times 3$ to denote the matrix representation of the cross product with the vector $v$, i.e.,
such that $[v]_{\times}w=v\times w$ for any $w\in 3$.
Proposition 2
Let $\nu_{\mathcal{F}}(t)$ represent a parametric curve. Then
$\dot{H}=H[-2M\dot{\nu}_{\mathcal{F}}]_{\times}\!,$ (6)
where the matrix $M\in 3\times 3$ is given by
$M=[u]_{\times}\frac{\left(I-uu^{\mathrm{T}}\right)\left(I-\nu_{\mathcal{F}}\nu_{\mathcal{F}}^{\mathrm{T}}\right)}{\lVert u^{\prime}\rVert\lVert\nu_{\mathcal{F}}\rVert}.$ (7)
Proof:
See Appendix -B. ∎
II-A3 Alternating Directions Method of Multipliers (ADMM)
@@
The basic idea of the ADMM-based solver introduced in [6] is to separate the constraints from the objective function using a different set of variables $\mathbf{z}$, and then solve separately in (1). More specifically, we rewrite the constraint $\mathbf{q}\in\varOmega$ using an indicator function $\varTheta$, and include it in the objective function (details can be found in [6]). We allow $\mathbf{z}=D(\mathbf{q})$ to replicate an arbitrary function of the main variables $\mathbf{q}$ (instead of being an exact copy in general ADMM formulation), to transform constraint $\mathbf{q}\in\Omega$ to $D(\mathbf{q})\in\mathcal{Z}$ to allow for an easier projection step in Equation 9b. In summary, we transform Equation 1 into
$\max\quad\varPhi(\mathbf{q})+\varTheta(\mathbf{z})$ (8)
$s.t.&D(\mathbf{q})-\mathbf{z}=0\end{array}$
@@
where $D(\mathbf{q})=[D_{1}(\mathbf{q})^{T},\dots,D_{l}(\mathbf{q})^{T}]^{T}$ is a vertical concatenation of different functions for different constraints. This makes each constraint set $\mathcal{Z}_{i}$ independent and thus can be computed separately in later updating steps. $D_{i}(\mathbf{q})$ is chosen that the new constraint set $\mathcal{Z}_{i}$ becomes simple to compute which is illustrated in later secitons.
@@
Checking the primal residuals alongside with the dual residuals
@@
after each iteration, the steps are reiterated until convergence when the primal and dual residuals are small, or divergence when primal and dual residual remains large after a fixed number of iterations.
We now provide the functions $D(\mathbf{q})$, the sets $\mathcal{Z}$, and the corresponding projection operators $\Pi_{\mathcal{Z}}$ for security constraints including co-observation security constraints (Section II-B), and reachability constraints (Section II-E-Equation 51). The latter are based on the definition of ellipse-region-constraint (Section II-C). Formulation of traditional path planning constraints like velocity constraints, convex obstacle constraints can be found in [6], thus are omitted in this paper.

### v2

@@ Abstract
We address security challenges in multi-robot systems (MRS) where adversaries may take control of compromised robots to gain unauthorized physical access to forbidden regions. We propose a novel multi-robot optimal planning algorithm that integrates mutual observations and introduces reachability constraints and co-observations between robots to enhance security. Our formulation can guarantee that, even with adversarial movements, compromised robots cannot breach forbidden regions without missing scheduled co-observations. We use ellipsoidal over-approximations to reachability regions for efficient intersection checking and gradient computation. Furthermore, to enhance system resilience and tackle feasibility challenges, we introduce the notion of sub-teams; these cohesive units replace individual robots along each trajectory; we plan cross-trajectory co-observation that use redundant robots to switch between different trajectories, securing multiple sub-teams without requiring modifications to the plans. We formulate cross-trajectories plans by solving a network-flow vertex path coverage problem on the checkpoint graph generated from the original unsecured MRS trajectories, providing the same security guarantees against plan-deviation attacks. We demonstrate the effectiveness and robustness of our proposed algorithm through simulations, which significantly strengthen the security of multi-robot systems in the face of adversarial threats.
@@ I Introduction
Multi-robot systems (MRS) have found wide applications in various fields. Although offering numerous advantages, the distributed nature and dependence on network communication make the MRS vulnerable to cyber threats, such as unauthorized access, malicious attacks, and data manipulation [1]. This paper addresses a specific scenario in which robots are compromised by physically masquerading attackers and directed toward forbidden regions, which may contain security-sensitive equipment or human workers. A countermeasure to these plan-deviation attacks [2, 3, 4] is to use the onboard sensing capabilities of the robots themselves to perform inter-robot co-observations and detect unusual behavior. These mutual observations establish a co-observation schedule alongside the planned path, ensuring that any attempts by a compromised robot to violate safety constraints (such as transgressing into forbidden regions) would break the observation plan and be promptly detected.
A preliminary version of this approach is presented in [5, 6]. That work extended the grid-world solution from [2] to continuous configuration space, and incorporated the co-observation planning as constraints in the alternating direction method of multipliers (ADMM)-based trajectory optimization solver to accommodate for more flexible objectives. However, such plans are not guaranteed to exist, and the secured trajectory always comes at the cost of overall system performance. As an extension of the previous works [5, 6], we introduce reachability analysis and network flow to the security problem to find a balance between security and performance.
@@
In this paper, we incorporate additional reachability analysis during the planning phase, incorporating constraints based on sets of locations that agents could potentially reach, referred to as reachability regions, as a novel perspective to the existing literature. This approach enforces an empty intersection between forbidden regions and the reachability region during trajectory optimization, preventing undetected attacks if an adversary gains control of the robots. We utilize an ellipsoidal boundary to constrain the search space and formulate the ellipsoid as the reachability region constraint. We present a mathematical formulation of the reachability regions as spatio-temporal constraints compatible with the solver in [5].
@@
Furthermore, we address the feasibility and optimality challenges. For a MRS with unsecured optimal trajectories (without security constraints), we introduce redundant robots and form them into sub-teams with the original ones. These redundant robots are assigned to establish additional co-observations, termed cross-trajectory co-observations, within and across different sub-teams. The proposed algorithm focuses on the movement plans for the additional robots, distinct from those dedicated to task objectives, indicating when they should stay with their current sub-team and when they should deviate to join another. This strategy allows sub-teams to preserve the optimal unsecured trajectories (as illustrated in Fig. 1) without requiring the entire sub-team to maintain close proximity to other teams during co-observation events. The cross-trajectory co-observation problem is transformed as a Multi-Agent Pathfinding (MAPF) problem on roadmap (represented as directed graph) and solved as a network flow problem [7].
@@
Figure 1: (1a) The attacker deviates the red robot into a forbidden region.(1b) Red and blue robots are scheduled to co-observe each other during the task, resulting in a new secured trajectory (solid lines) replacing the optimal ones (dashed line). (1c) The blue team sends one robot to observe the red team (solid blue line) while having the rest of the robots follow the optimal trajectory.
@@
Related research. The trajectory planning problem of MRS remains as a subject of intense study for many decades, and optimization is a common approach in such areas. Optimization-based approaches are customizable to a variety of constraints (e.g. speed limit, avoiding obstacles) and task specifications (e.g. maximum surveillance coverage, minimal energy cost). In contrast to our work in Section II, many motion planning tasks require a non-convex constraint problem formulation, most contributors focused on convex problems, and only allow for a few types of pre-specified non-convex constraints through convexification[8][9][10]. Several optimization techniques like mixed integer programs with quadratic terms (MIQP) [11] and ADMM [12] have been used to reduce computational complexity and to incorporate more complex non-convex constraints.
@@
Graph-based search is another extensively explored approach in trajectory planning problems, such as formation control [13, 14] and MAPF problems [15]. In traditional MAPF formulations, environments are abstracted as graphs, with nodes representing positions and edges denoting possible transitions between positions, and solved as a network flow problem [7, 16]. This formulation allows for the application of combinatorial network flow algorithms and linear program techniques, offering efficient and more flexible solutions to the planning problem.
Reachability analysis is essential for security and safety verification in cyber-physical systems (CPSs) [17, 18], often involving an over-approximation of reachable space to verify safety properties. Geometric methods like zonotopes and ellipsoids are commonly used to enclose reachability sets compactly [19, 20, 21]. For online safety assessments against cyber attacks, [22] compute reachable CPS states under attacks and compared them with a safe region based on state estimation. By incorporating an additional security measure that provides two secured states, the reachability region in our work is over-approximated using ellipsoids. This is inspired by the ellipsoidal heuristic sampling domain in [23] for the $\mathtt{RRT^{*}}$ algorithm to simplify the sampling region between start and goal locations. Ellipsoids are also used in other path-planning methods like Iterative Regional Inflation by Semidefinite programming (IRIS) [24, 25] and the Safe Flight Corridor (SFC) [26, 27]. Differing from the reachability region that requires mapping all reachable states given several known states, the ellipsoids of IRIS and SFC focus on approximating the safe collision-free space rather than addressing security applications.
Paper contributions. Two main contributions have been presented.
@@
  • We present an innovative method to integrate reachability analysis into the ADMM-based optimal trajectory solver for multi-robot systems, preventing attackers from executing undetected attacks by simultaneously entering forbidden regions and adhering to co-observation schedules.
@@
  • We introduce additional robots to form sub-teams for both intra-sub-team and cross-sub-team co-observations. A new co-observation planning algorithm is formulated that can generate a resilient multi-robot trajectory with a co-observation plan that still preserves the optimal performance against arbitrary tasks. We also find the minimum redundant robots required for the security.
@@
The rest of this paper is organized as follows. Section II introduces the ADMM-based optimal trajectory solver and the security constraints. Section III introduces the enhanced security planning algorithm through cross-trajectory co-observation. Section IV concludes this article.
Notation. In this paper, we use non-bold symbols to denote single-agent states (e.g. $q_{ij}$) and scalars, and bold symbols to represent aggregated states of single robot and multiple robots (e.g. $\mathbf{q}$).
@@ II Secured Multi-robot trajectory planning
We formulate the planning problem as an optimal trajectory optimization problem to minimize arbitrary smooth objective functions. We denote the trajectory as $\mathbf{q}_{i}=[q_{i0}\dots q_{iT}]\in m\times T$, where $q_{ij}\in m$ is the waypoint of agent $i$ in a $m$ dimensional state space, $T$ is the time horizon. For a total of $n_{p}$ robots, trajectories can be represented as an aggregated vector $\mathbf{q}=\operatorname{stack}(\mathbf{q}_{1},\dots,\mathbf{q}_{n_{p}})\in mn_{p}\times T$, where $\operatorname{stack}(\cdot)$ denotes the vertical stacking operation. The overall goal is to minimize or maximize an objective function $\varPhi(\mathbf{q}):nmT\to$ under a set of nonlinear constraints described by a set $\Omega$, which is given by the intersection of spatio-temporal sets given by traditional path planning constraints and the security constraints (co-observation schedule, reachability analysis). Formally:
@@
Robots are tasked with navigating an unknown task space to collect sensory data and reconstruct a field (as illustrated in Fig. 2a). The task space is modeled as a grid, where each grid point has an associated value tracked by a Kalman Filter (KF) [28]. The KF estimates uncertainty at each point through its covariance $P_{j}$ updated based on the measurements taken by robots along the trajectory $\mathbf{q}$. Measurement quality, modeled by a Gaussian radial basis function, decreases with distance from the robot. The optimization objective $\varPhi(\mathbf{q})=\max_{j}P_{j}(\mathbf{q})$ is to minimize the maximum uncertainty in reconstructing the field (detailed in [5]).
(a) Map exploration task
(b) Continuous world trajectory
Figure 2: (2a) Illustration of a 3 robot map exploration task case. (2b) The unsecured trajectory design optimized for a map exploration task. Potential security breaches, indicated by red dashed lines, highlight paths that could allow unauthorized access to forbidden regions.
Ideally, robots are expected to spread out in a Boustrophedon pattern to efficiently explore the task space, optimize coverage, and minimize reconstruction uncertainty (Fig. 2b). However, we consider the possibility of threats following the physical masquerade attack model [2], where a compromised insider (robot) masquerading as a properly functioning agent, attempts to gain access to unauthorized locations without detection. The attacker leverages full knowledge of the motion plan and exploits the compromised robot to provide false self-reports to the central entity (CE). These malicious deviations remain undetected as long as observations from uncompromised robots align with the motion plan. The corresponding security requirement can then be formally defined as:
@@
A multi-robot trajectory plan is secured against plan-deviation attacks if it ensures that any potential deviations to these forbidden regions will cause the corresponding robot to miss their next co-observation with other robots.
@@ II-A1 Differentials
We define the differential of a map $f(x):d_{1}\to d_{2}$ at a point $x_{0}$ as the unique matrix $\partial_{x}f\in d_{1}\times d_{2}$ such that
$\left.\frac{\mathrm{d}}{\mathrm{d}t}f\bigl{(}x(t)\bigr{)}\right|_{t=0}=\partial_{x}f\bigl{(}x(0)\bigr{)}\dot{x}(0),$ (2)
@@
where $t\mapsto x(t)\in d_{1}$ is a smooth parametric curve such that $x(0)=x$ with any arbitrary tangent $\dot{x}(0)$. For brevity, we will use $\dot{f}$ for $\frac{\mathrm{d}}{\mathrm{d}t}f$ and $\partial_{x}f$ for $\frac{\partial f}{\partial x}$.
@@
With a slight abuse of notation, we use the same notation $\partial_{x}f$ for the differential of a matrix-valued function with scalar arguments $f:d_{1}\times d_{3}\to d_{2}$. Note that in this case (2) is still formally correct, although the RHS needs to be interpreted as applying $\partial_{x}f$ as a linear operator to $\dot{x}$.
II-A2 Alternating Directions Method of Multipliers (ADMM)
@@
The basic idea of the ADMM-based solver introduced in [5] is to separate the constraints from the objective function using a different set of variables $\mathbf{z}$, then solve separately in (1). More specifically, we rewrite the constraint $\mathbf{q}\in\varOmega$ using an indicator function $\varTheta$ and include it in the objective function (details can be found in [5]). We allow $\mathbf{z}=D(\mathbf{q})$ to replicate an arbitrary function of the main variables $\mathbf{q}$ (instead of being an exact copy in general ADMM formulation), to transform constraint $\mathbf{q}\in\Omega$ to $D(\mathbf{q})\in\mathcal{Z}$ to allow for an easier projection step in Eq. 4b. In summary, we transform Eq. 1 into
$\max\quad\varPhi(\mathbf{q})+\varTheta(\mathbf{z}),$ (3)
[…]
@@
where $D(\mathbf{q})=[D_{1}(\mathbf{q})^{T},\dots,D_{l}(\mathbf{q})^{T}]^{T}$ is a vertical concatenation of different functions for different constraints. This makes each constraint set $\mathcal{Z}_{i}$ independent and thus can be computed separately in later updating steps. $D_{i}(\mathbf{q})$ is chosen such that the new constraint set $\mathcal{Z}_{i}$ becomes simple to compute which is illustrated in later sections.
@@
Checking the primal residuals alongside the dual residuals
@@
after each iteration, the steps are reiterated until convergence when the primal and dual residuals are small, or divergence when primal and dual residuals remain large after a fixed number of iterations.
Remark 1
In practical application, (4a) is solved iteratively via a nonlinear optimization solver such as fmincon [30]. The use of explicit gradients of the constraint functions (derived below) greatly enhances the solver’s efficiency and accuracy (with respect to the use of numerical differentiation). At the same time, (4b) and (4c) have closed form solutions.
We now provide the functions $D(\mathbf{q})$, the sets $\mathcal{Z}$, and the corresponding projection operators $\Pi_{\mathcal{Z}}$ for security constraints including co-observation security constraints (Section II-B), and reachability constraints (Section II-E-Eq. 30). The latter is based on the definition of ellipse-region-constraint (Section II-C). Formulation of traditional path planning constraints like velocity constraints and convex obstacle constraints can be found in [5], thus are omitted in this paper.